deathscarewolf:

The sets of numbers 6, 8, 10 and 5, 12, 13 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

jhonyy9:

ok so just and example 3,4 and 5 are Pythagorean triples bc. 5^2 = 3^2 +4^2 25 = 9 +16 25 = 25 hope helped